OpenStudy (anonymous):

In a sample of 26 hand-held calculators, 20 are known to be nonfunctional. If 6 of these calculators are selected at random, what is the probability that exactly 4 in the selection are nonfunctional? Round to the nearest thousandth. A. 0.667 B. 0.316 C. 0.769 D. 0.300 E. 0

OpenStudy (amistre64):

20 out of 26 are bad.... thats a ratio of: 20/26 = 10/13 = 1/1.3 1:1.3 = 6:x i think will give the probablity of how many are bad

OpenStudy (amistre64):

then 4/6 should give you something along the lines of figuring out the rest...